OpenStudy (anonymous):

(ill fan anyone who can answer this) Which phrase represents the algebraic expression three over x plus five? five divided by a number, increased by three five greater than a number divided into three the sum of three and a number divided by five five more than the quotient of a number and three

OpenStudy (anonymous):

The original expression is a little ambiguous, it could mean: \[\frac{ 3 }{ x+5 } or \frac{ 3 }{ x } + 5\]Which means the answer could be the second choice or the fourth choice

OpenStudy (anonymous):

o.o that's seems 50/50

OpenStudy (anonymous):

The second and fourth options seem equivalent to me. They both describe this: \[\frac{ 3 }{ x } + 5\]I would go with the second choice. It's more specific, and it's accurate

OpenStudy (anonymous):

thats what I went for